OpenStudy (goformit100):

Prove that the fraction (21n+4)/(14n+3) is irreducible for every natural number n.

OpenStudy (goformit100):

@mikaela19900630

OpenStudy (goformit100):

@qweqwe123123123123111

OpenStudy (amistre64):

when does the top equal the bottom?

OpenStudy (goformit100):

I am unable to understand your question.

OpenStudy (amistre64):

since the top and bottom of the fraction are linear and do not cancel, then we would want to know when the top equal the bottom in order that they could cancel each other out

OpenStudy (goformit100):

is it 0 ?

OpenStudy (amistre64):

lets try (21n+4)/(14n+3) (21(0)+4)/(14(0)+3) 4/3 nope

OpenStudy (goformit100):

ok then there is no natural number of such type.

OpenStudy (amistre64):

correct, but that is not a "proof"

OpenStudy (goformit100):

ok Thankyou sir I got it :)

OpenStudy (amistre64):

21n + 4 = 14n + 3 7n = -1 n = -1/7 is the only value of n which would make this reduce to 1. but -1/7 is not a natural number

OpenStudy (goformit100):

ok

OpenStudy (goformit100):

Thanks Again MOD

OpenStudy (experimentx):

21n + 4 = (14n + 3)k show that 'k' is not natural number,

OpenStudy (goformit100):

ok

OpenStudy (goformit100):

@experimentX thankyou sir

OpenStudy (experimentx):

try making use of induction.

OpenStudy (goformit100):

Sure sir

OpenStudy (goformit100):

oh yeah I proved it :)